OpenStudy (anonymous):

Please Help. How do you find the total words of a book without counting each one of them? So not like an exact exact number since I am surely not going to count all the words.

OpenStudy (anonymous):

Count the number on a page and multiply by the number of pages.

OpenStudy (anonymous):

okaii thnx :D

OpenStudy (unklerhaukus):

words per book = words per line x lines per page x pages per book \[\frac{\text {words}}{\text{book}} = \frac{\text{words}}{\text{line}} \times \frac{\text{lines}}{\text{page}}\times \frac{\text{pages}}{\text{book}}\]